OpenStudy (anonymous):

how do you simplify /factor this.....??????? HELP 2x^2-x-3, to find the limit of 2x^2-x-3/x+1

OpenStudy (anonymous):

2x^2-x-3 = 2x^2-3x+2x-3

OpenStudy (anonymous):

(x^2+?)(x-?)

OpenStudy (anonymous):

2x^2-x-3 2x^2-3x+2x-3 x(2x-3)+1(2x-3) (2x-3)(x+1)

OpenStudy (anonymous):

thank you @telltoamit

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

do u know how do factor x^3+8?

OpenStudy (anonymous):

umm, no idea, sorry

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

wait, i got it: x3 + 8 = x3 + 23 = (x + 2)[x2 - (x)(2) + 22] = (x + 2) (x2 - 2x + 4)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yvw